X=4i-2j

Y=i+5j

determine; i. the magnitude of x+y

ii. x-4y

the unit vector y^

1
Expert's answer
2021-07-02T11:42:22-0400

Given:

"x=4i-2j"

"y=i+5j"

(i)

"|x+y|=|4i-2j+i+5j|\\=|5i+3j|=\sqrt{5^2+3^2}=\sqrt{34}=5.8"

(ii)

"|x-4y|=|4i-2j-4i-20j|\\=|-22j|=22"